# SpikeDet: Better Firing Patterns for Accurate and Energy-Efficient Object Detection with Spiking Neuron Networks

### v3

@@ B-2 More Implementation Details
The MDSNet structure is detailed in Table VIII. For SpikeDet-L, we increase the channel dimensions of MDSNet104 by a factor of 1.25 to improve model performance. On the Gen1 dataset, we employ the zoom-in and zoom-out augmentation strategies from [65]. The model is trained for 100 epochs with a batch size of 64. On the URPC 2019 dataset, we use mosaic augmentation and, similar to [47, 28, 68], resize images to 320$\times$320 and train for 300 epochs with a batch size of 64.
@@ Appendix C More Visualization
In this section, we present the visualization results of our model on the URPC 2019 dataset, as shown in Fig. 9. These results demonstrate that our model can achieve excellent performance on the object detection sub-tasks.
Figure 9: Detection result on the URPC 2019 dataset.

### v4

@@ B-2 More Implementation Details
The MDSNet structure is detailed in Table X. For SpikeDet-L, we increase the channel dimensions of MDSNet104 by a factor of 1.25 to improve model performance. On the Gen1 dataset, we employ the zoom-in and zoom-out augmentation strategies from [gehrig_recurrent_2023]. The model is trained for 100 epochs with a batch size of 64. For URPC 2019, ExDARK, and CrowdHuman datasets, we employ mosaic augmentation. Specifically, following [li_suyolo_2025, wang_ulo_2022, pan_optimization_2024], we resize images to $\times320$ for URPC 2019, while using $\times640$ for the other datasets. All models are trained for 300 epochs with a batch size of 64.
@@ Appendix C More Visualization
In this section, we present the visualization results of our model on the URPC 2019, ExDARK, and CrowdHuman datasets, as shown in Figs. 12, 13, and 14. These results demonstrate that our model achieves excellent performance on the object detection subtasks.